# Federated Learning with Dynamic Client Arrival and Departure: Convergence and Rapid Adaptation via Initial Model Construction

### v2

@@ C.1 Cost Analysis with Variable Clients
  • Storage: One gradient of size $W$ is stored for session $s$.
Coefficient Calculation (for session $m$, $m\in\{P+1,\dots,S\}$, at $g=mT$):
  • Computation: Computing softmax-based coefficients over the $(m-P)$ stored items incurs
    $\mathcal{O}(m-P)$
    cost for vector norms and exponentials.
Total Persistent Storage:
  • Pilot model: One model of size $W$ (intermediate checkpoints discarded).
  • Per post-pilot session $s$:
    • Final model checkpoint: $W$,
    • Computed gradient: $W$.
  • Overall growth: Linear increase of $2W$ per post-pilot session, scaling with $(S-P)$.
C.2 Storage Overhead
Persistent storage grows linearly with $S-P$ (one model and one gradient per session), which remains practical.
C.3 Justification for Computation and Communication Requirements
The upfront cost of pilot training and gradient computation is offset by accelerated adaptation to shifting client distributions, reducing the total number of global rounds (e.g., avoiding $50+$ recovery rounds required by standard baselines).
C.4 Mitigating Storage for Matrices $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$
To bound storage growth, one may:
  • Selective retention: Store models/gradients that maximize distributional diversity, preserving initialization quality with fewer checkpoints.
Appendix D Experimental Setup, Reproducibility, and Computational Resources
D.1 Setup and Reproducibility
@@
How hyperparameters were chosen: We selected all hyperparameters such that the training within each session converges by its final round. This ensures a fair comparison across different algorithms and datasets.
Baseline variants: AG News uses the average and previous baselines. All other datasets use average, previous, and continuous baselines.
20 Clients Configuration
  • num_sessions=6
  • acg_beta=0.01, acg_lambda=0.85
  • prox_alpha=1.0 for FedProx
  • Algorithms:
    • All datasets: FedAvg, FedProx, SCAFFOLD, FedACG
    • TinyImageNet: MOON (moon_mu=1.0, moon_tau=1.0), FedProx, FedACG
  • kl_coefficient = 1.0 for all datasets except AG News (0.5)
  • similarity_scale = 10.0 for all datasets except AG News (800)
  • Random seeds: 0, 1, 2 (TinyImageNet uses 100, 200, 300)
@@ D.2 Compute Resources
LABEL:table:resource_20_appendix presents a comprehensive overview of average CPU time and maximum memory usage for our experiments with 20 clients. The average is taken over three random seeds.
Table 15: Average CPU time and average maximum memory usage across three random seeds for our experiments with 20 clients.
@@
LABEL:table:resource_100_appendix presents a comprehensive overview of average CPU time and average maximum memory usage for our experiments with 100 clients. The average is taken over three random seeds.
Table 16: Average CPU time and average maximum memory usage across three random seeds for our experiments with 100 clients.
@@
Appendix E Proof of Theorem 1
Using the $\beta$-smoothness of the global loss function (Assumption 1), we have
Plugging (6) and (7) into (9) and taking the expectation give
$\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]$ (10)
We first bound term (a) as follows.
where (i) we have used the Cauchy-Schwarz inequality $\|\bm{a}+\bm{b}\|^{2}\leq 2\|\bm{a}\|^{2}+2\|\bm{b}\|^{2}$, (ii) we used the convex property of 2-norm, (iii) we used Eq (3.5) in chapter 3 of [51] and the definition of variance of the gradients $(\sigma_{k}^{(g),e-1})^{2}$ evaluated at the particular local step $e-1$ for the parameter realization $\bm{w}_{k}^{(g),e-1}$. Expanding $(\sigma_{k}^{(g),e-1})^{2}$ yields
$(\sigma_{k}^{(g),e-1})^{2}$
where $d$ denotes the feature vector of data point $d$, and $\bm{\lambda}_{k}^{(g)}$ and $(\widetilde{\sigma}_{k}^{(g)})^{2}$ denote the mean and sample variance of data points in $\mathcal{D}_{k}^{(g)}$, which are gradient independent. Also, in (i) of (E), we used the Cauchy-Schwarz inequality, and in (ii), we used the fact that $\sum_{d\in\mathcal{D}_{k}^{(g)}}\left(d-\bm{\lambda}_{k}^{(g)}\right)=0.$ Plugging (E) into (E) gives
Using this, (E) becomes
Rearranging the terms, we have
$\mathbb{E}_{g}\big{[}F^{(g)}(\bm{w}^{(g+1)})\big{]}$
Term (b) in (E) can be upper bounded as
$\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
The last term in (18) can be further upper bounded
Combining (19) with (18) gives
$\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
Taking the summation of both sides in (20) across local training steps
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
After some simplification, we obtain
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
Assuming
$\eta^{(g)}\leq\Big{(}2\beta^{2}\sqrt{e_{k}^{(g)}(e_{k}^{(g)}-1)}\Big{)}^{-1},$
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
Plugging (23) into (E) gives
To eliminate term (c) in (24), we need to make
we require
Also, we aim to make
$01-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)>$ 0 . (28)
Thus, we get the condition:
Assume there exists a set of constants $\Lambda^{(g)}$, $\forall g$, where:
Then we will have
and
$\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}$
$\frac{1}{G}\sum_{g=1}^{G}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}$

### v3

@@ VI-5 Local Model Computation
[…]
Quantification of Performance Gain: We quantify performance based on the model’s ability to rapidly adapt with two metrics: (1) Time-to-Accuracy and (2) Accumulated Accuracy Gain. Time-to-Accuracy (Adaptation Speed) metric measures the latency of adaptation following the start of a new session. It is defined as the number of global rounds $t$ (within the current session $s$) required for an algorithm to recover to a specific percentage ($\rho$) of the peak accuracy achieved by the proposed scheme in that session. Let $Acc_{s}^{(\text{alg})}(t)$ denote the accuracy of an algorithm at the $t$-th global round of session $s$, and let $Acc^{*}_{s,\text{prop}}$ be the maximum accuracy achieved by the proposed method within session $s$. The Time-to-Accuracy $T_{\rho}$ is defined as:
$T_{\rho}=\min\{t\in[1,T]\mid Acc_{s}^{(\text{alg})}(t)\geq\rho\cdot Acc^{*}_{s,\text{prop}}\}$ (31)
where $T$ is the total number of global rounds per session, and $\rho\in\{0.95,0.97\}$. A lower $T_{\rho}$ indicates that the algorithm requires fewer global rounds (and consequently less latency and energy) to adapt to the new distribution. If the target is not reached within the session, $T_{\rho}$ is denoted as $\infty$.
While Time-to-Accuracy measures convergence speed to a specific target, it fails to capture the stability or sustained magnitude of performance throughout the session. To address this, we define the Accumulated Accuracy Gain, which quantifies the cumulative performance advantage of the proposed method over a baseline.
Calculated as the sum of accuracy differentials between the proposed method and a baseline $b$, the total gain $G_{\text{total}}$ for session $s$ is defined as:
$G_{\text{total}}=\sum_{t=1}^{T}\left(Acc_{s}^{(\text{prop})}(t)-Acc_{s}^{(b)}(t)\right)$ (32)
A positive $G_{\text{total}}$ indicates that the proposed method maintains a superior accuracy trajectory compared to the baseline across the session’s global rounds.
Table VII presents the quantitative results for Time-to-Accuracy and Accumulated Accuracy Gain, highlighting the efficiency of the proposed dynamic initialization strategy. The most significant advantage is the substantial reduction in latency and energy consumption, computed using our communication and computation model. Across all scenarios, the proposed method demonstrates rapid adaptation, consistently recovering the target accuracy in just a single round. In the challenging TinyImageNet tasks (specifically under the SCAFFOLD algorithm), this efficiency contrasts sharply with the Continuous baseline, which requires 150 rounds to reach the same threshold. This convergence gap translates into substantial resource savings; the proposed method consumes only 1.43 kJ of energy compared to the 214.33 kJ required by the Continuous baseline, representing a $150\times$ reduction in resource overhead.
Furthermore, the Accumulated Accuracy Gain column confirms that the proposed method maintains a sustained performance advantage throughout the entire adaptation phase. This improvement is most pronounced in scenarios with zero overlap (e.g., SCAFFOLD on TinyImageNet), where our method yields a total accuracy gain of $867.1\%$ over the Continuous baseline. Against the Previous baseline, our method consistently achieves positive total gains, validating that the proposed weighted initialization strategy provides a superior starting point compared to simply reusing the last session’s model.
TABLE VII: Comparison of Time-to-Accuracy (95% Target for TinyImageNet, 97% Target for SVHN) and Accumulated Accuracy Gain.
Dataset Overlap Dirichlet $\bm{\alpha}$ FL Alg. Method Time-to-Accuracy (95%/97% Target) Accumulated
Rounds Latency (s) Energy (kJ) Speedup Accuracy Gain (%)
TinyImageNet 0.2 0.7 FedACG Proposed 1 113.6 1.43
Previous (Baseline) 11 1,249.4 15.72 $11.0\times$ +139.6+139.6
Continuous (Baseline) 55 6,247.2 78.59 $55.0\times$ +542.4+542.4
Average (Baseline) 7 795.1 10.00 $7.0\times$ +61.1+61.1
TinyImageNet 0.0 0.3 SCAFFOLD Proposed 1 113.6 1.43
Previous (Baseline) 14 1,590.2 20.00 $14.0\times$ +85.7+85.7
Continuous (Baseline) 150 17,037.7 214.33 $150.0\times$ +867.1+867.1
Average (Baseline) 28 3,180.4 40.01 $28.0\times$ +243.5+243.5
SVHN 0.0 0.7 MOON Proposed 1 6.9 0.08
Previous (Baseline) 5 34.4 0.39 $5.0\times$ +11.5+11.5
Continuous (Baseline) 52 358.1 4.04 $52.0\times$ +395.1+395.1
Average (Baseline) 14 96.4 1.09 $4.0\times +11$6.1+116.1
SVHN 0.2 0.3 FedProx Proposed 1 6.9 0.08
Previous (Baseline) 4 27.5 0.31 $4.0\times$ +8.0+8.0
Continuous (Baseline) 70 482.1 5.44 $70.0\times$ +517.6+517.6
Average (Baseline) 10 68.9 0.78 $0.0\times +11$5.0+115.0
Figure 4: Impact of system bandwidth on Total Latency (left) and Total Energy Consumption (right) required to reach 97% of the proposed scheme’s peak accuracy on the SVHN dataset. The proposed method consistently achieves the lowest latency and energy consumption across all bandwidth configurations (50–200 MHz). Insets highlight the performance gap between the Proposed and Previous schemes at a finer scale.
Impact of System Bandwidth: We further investigate the robustness of the proposed framework to communication resource constraints by varying the system bandwidth from 50 MHz to 200 MHz. Figure 4 illustrates the total latency and energy consumption required to recover 97% of the maximum accuracy for our proposed scheme on the SVHN dataset. The results demonstrate that our method consistently achieves the target accuracy with fewer communication rounds. Consequently, our approach effectively saves communication costs and yields the lowest resource overhead across all bandwidth configurations. While higher bandwidth naturally reduces transmission time for all methods, the proposed algorithm maintains a distinct efficiency advantage as highlighted by the insets comparing it to Previous baseline. This confirms that the efficiency gains are driven by the reduction in required training rounds and algorithmic convergence rather than being dependent on specific channel conditions.
Other Experiments: Our online technical report [chang2025federatedlearningdynamicclient] provides extensive supplementary results for our other benchmark datasets (MNIST, Fashion-MNIST, SVHN, CIFAR-10, and CIFAR-100) across systems with varying devices. These additional experiments demonstrate the robustness of our approach under several alternative non-IID data distributions as well.
VII Conclusion
In this paper, we addressed the structural challenges inherent to Federated Learning in wireless edge networks, specifically those driven by dynamic device arrival and departure. We presented a theoretical convergence analysis that explicitly accounts for the shifting optimization goals associated with evolving device sets, quantifying the impact of gradient noise and data heterogeneity. Guided by these theoretical insights, we proposed a plug-and-play initialization strategy that accelerates adaptation by constructing a weighted average of historical models based on gradient similarity. This approach effectively bridges distributional shifts between sessions without requiring server-side knowledge of future device activities. Extensive simulations demonstrate that our method achieves convergence speedups typically an order of magnitude or more compared to baselines and substantial energy savings compared to state-of-the-art baselines. Future work will focus on developing adaptive triggering mechanisms to determine when a new session initialization is required, further minimizing computational overhead in highly volatile environments.
Appendix A Communication and Computation Modeling Details
To evaluate the real-world performance of Algorithm 1 in a realistic wireless edge network, we model the latency and energy constraints imposed by the physical environment. We consider a Time Division Duplex (TDD) system to separate uplink and downlink transmission intervals. Within each interval, we assume an Orthogonal Frequency Division Multiple Access (OFDMA) scheme where devices are assigned orthogonal resource blocks, ensuring an interference-free environment.
A-A Channel Modeling
The wireless channel between device $k$ and the server is characterized by both large-scale and small-scale fading. The total channel gain $g_{k}[n]$ at time slot $n$ is given by:
$|g_{k}[n]|^{2}=\psi_{k}\cdot|h_{k}[n]|^{2},$ (33)
where $\psi_{k}$ denotes the large-scale fading coefficient and $h_{k}[n]$ represents the small-scale fading component.
A-A1 Large-Scale Fading
We adopt the 3GPP Urban Micro (UMi) Street Canyon model defined in the technical report TR 38.901 [3gpp38901], the standard for 5G/6G simulations in dense urban environments. The coefficient $\psi_{k}$ captures the combined impact of path loss and shadowing. Shadowing is modeled via a random variable $\xi$ following a zero-mean Gaussian distribution, $\xi\sim\mathcal{N}(0,\sigma_{SF}^{2})$. For the UMi-LoS scenario, the standard deviation is $\sigma_{SF}=4\text{ dB}$. The linear large-scale fading coefficient $\psi_{k}$ is obtained by converting the total path loss from the logarithmic scale:
$\psi_{k}=10^{-(PL_{\text{UMi-LoS}}+\xi)/10}.$ (34)
We assume $\psi_{k}$ remains constant during a single communication round. The path loss $PL_{\text{UMi-LoS}}$ is divided into two segments based on a “Break Point” distance ($d^{\prime}_{BP}$):
$PL_{\text{UMi-LoS}}[\text{dB}]=\begin{cases}PL_{1},&\text{if }10\text{m}\leq d_{\text{2D}}\leq d^{\prime}_{BP}\\ PL_{2},&\text{if }d^{\prime}_{BP}<d_{\text{2D}}\leq 5\text{km}\end{cases}$ (35)
where $d_{\text{2D}}$ is the 2D distance between the device and the Base Station (BS). The segment path losses are:
$PL_{1}$ $=32.4+21\log_{10}(d_{\text{3D}})+20\log_{10}(f_{c}),$ (36)
$PL_{2}$ $=32.4+40\log_{10}(d_{\text{3D}})+20\log_{10}(f_{c})$
$\quad-9.5\log_{10}\left((d^{\prime}_{BP})^{2}+(h_{\text{BS}}-h_{\text{UT}})^{2}\right),$ (37)
where $f_{c}$ is the carrier frequency in GHz, $h_{\text{BS}}$ is the BS height, $h_{\text{UT}}$ is the user terminal height, and $d_{\text{3D}}=\sqrt{d_{\text{2D}}^{2}+(h_{\text{BS}}-h_{\text{UT}})^{2}}$ is the 3D distance.
Following TR 38.901 [3gpp38901], the break point distance is calculated using effective antenna heights to account for the environment height ($h_{E}=1.0$ m). The effective heights are $h^{\prime}_{\text{BS}}=h_{\text{BS}}-h_{E}$ and $h^{\prime}_{\text{UT}}=h_{\text{UT}}-h_{E}.$ Consequently, the break point distance is $d^{\prime}_{BP}=(4h^{\prime}_{\text{BS}}h^{\prime}_{\text{UT}}f_{c}\times 10^{9})/c,$ where $c=3\times 10^{8}$ m/s.
A-A2 Device Distribution
We consider a single-cell system where the server is at the center of a circular coverage area. Device locations are modeled as a Poisson Point Process (PPP) with uniform areal density within the cell, conditioned on a radius $R_{\text{cell}}=250$ m and a minimum distance $d_{\text{min}}=10$ m. The Euclidean 2D distance $d_{\text{2D},k}$ of device $k$ follows the probability density function (PDF):
$f(d)=\frac{2d}{R_{\text{cell}}^{2}-d_{\text{min}}^{2}},\quad d_{\text{min}}\leq d\leq R_{\text{cell}}.$ (38)
A-A3 Small-Scale Fading
The factor $h_{k}[n]$ is modeled as a first-order time-varying Gauss-Markov process:
$h_{k}[n]=\rho_{k}h_{k}[n-1]+\sqrt{1-\rho_{k}^{2}}e_{k}[n],$ (39)
where $e_{k}[n]\sim\mathcal{CN}(0,1)$ is an i.i.d. circularly symmetric complex Gaussian random variable. The temporal correlation coefficient $\rho_{k}$ is determined by the zeroth-order Bessel function of the first kind, $J_{0}(\cdot)$, as $\rho_{k}=J_{0}(2\pi f_{d}\tau)$, where $\tau$ is the slot duration and $f_{d}$ is the maximum Doppler frequency, given by $f_{d}=v_{k}f_{c}\times 10^{9}/c$. To ensure the channel remains approximately constant within a slot ($\tau\ll T_{c}$), we calculate the coherence time $T_{c}\approx 0.423/f_{d}$ and select $\tau=0.5$ ms.
A-B Instantaneous Data Rate and Latency
Based on the TDD assumption, the Uplink Instantaneous Data Rate (device $k$ to server) at time slot $n$ is:
$R_{k}^{\text{UL}}[n]=W\log_{2}\left(1+\frac{P_{k}^{\text{tx}}\psi_{k}|h_{k}[n]|^{2}}{N_{0}W}\right),$ (40)
where $W$ is the channel bandwidth per device, $P_{k}^{\text{tx}}$ is the transmission power of device $k$, and $N_{0}$ is the noise power spectral density. The Downlink Instantaneous Data Rate $R_{k}^{\text{DL}}[n]$ is calculated similarly using the server’s transmission power $P_{\text{server}}^{\text{tx}}$.
The total round-trip communication latency $T_{k}^{\text{comm}}$ for device $k$ to transmit model parameters of size $Z$ bits is:
$T_{k}^{\text{comm}}=T_{k}^{\text{DL}}+T_{k}^{\text{UL}},$ (41)
where $T_{k}^{\text{DL}}$ and $T_{k}^{\text{UL}}$ are the minimum durations satisfying $\sum_{n=1}^{T_{k}^{\text{DL}}/\tau}R_{k}^{\text{DL}}[n]\cdot\tau\geq Z$ and $\sum_{n=1}^{T_{k}^{\text{UL}}/\tau}R_{k}^{\text{UL}}[n]\cdot\tau\geq Z$, respectively.
A-C Energy Consumption and Computation
A-C1 Communication Energy
The total communication energy for device $k$ comprises the energy for receiving the global model and transmitting the local update:
$E_{k}^{\text{comm}}=P_{k}^{\text{rx}}\cdot T_{k}^{\text{DL}}+P_{k}^{\text{tx}}\cdot T_{k}^{\text{UL}},$ (42)
where $P_{k}^{\text{rx}}$ is the receiving circuit power. Typically $P_{k}^{\text{tx}}\gg P_{k}^{\text{rx}}$, making uplink transmission the dominant factor.
A-C2 Computation Latency and Energy
Let $N_{p}({\mathcal{M}})$ denote the number of parameters in model $\mathcal{M}$ and $\phi(\mathcal{M})$ the computational density ratio (FLOPs/param). To account for hardware capabilities, let $\chi$ denote the processor efficiency (FLOPs/cycle). Given that device $k$ performs $e_{k}^{(s,t)}$ local SGD iterations with mini-batch size $B^{(s,t)}$ during round $t$ of session $s$, the local computation time $T_{k}^{\text{comp}}$ is:
$T_{k}^{\text{comp}}=\frac{N_{p}(\mathcal{M})\cdot\phi(\mathcal{M})\cdot e_{k}^{(s,t)}\cdot B^{(s,t)}}{f_{k}\cdot\chi},$ (43)
where $f_{k}$ is the CPU clock frequency. The computation energy is modeled via the effective chipset capacitance $\xi_{k}$:
$E_{k}^{\text{comp}}=\xi_{k}\left(\frac{N_{p}(\mathcal{M})\cdot\phi(\mathcal{M})\cdot e_{k}^{(s,t)}\cdot B^{(s,t)}}{\chi}\right)(f_{k})^{2}.$ (44)
Appendix B Proof of Theorem 1
To streamline the derivation, we unify the session index $s\in\{1,\dots,S\}$ and the intra-session round index $t\in\{1,\dots,T\}$ from the main text into a single cumulative global round index $g\in\{1,\dots,G\}$, where $G=ST$. The mapping between these notations is given by g=(s1)T+tg=(s-1)T+t. For instance, in a system where $T=10$, the 5th round of the 3rd session ($s=3,t=5$) corresponds to the cumulative index g=(31)10+5=25g=(3-1)10+5=25.
Consequently, all parameters are mapped to this unified index $g$. For time-varying parameters such as the learning rate, we define $\eta^{(g)}\equiv\eta^{(s,t)}$. For parameters that remain constant throughout a session, such as the available device set $\mathcal{K}^{(s)}$, the notation $\mathcal{K}^{(g)}$ refers to the set associated with the unique session $s$ containing round $g$ (i.e., $\mathcal{K}^{(g)}\equiv\mathcal{K}^{(s)}$ holds for all $g$ such that $\lceil g/T\rceil=s$).
Using the $\beta$-smoothness of the global loss function Assumption (1), we have
Plugging (45) and (46) into (48) and taking the expectation give
$\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]$ (49)
We first bound term (a) as follows.
where (i) we have used the Cauchy-Schwarz inequality $\|\bm{a}+\bm{b}\|^{2}\leq 2\|\bm{a}\|^{2}+2\|\bm{b}\|^{2}$, (ii) we used the convex property of 2-norm, (iii) we used Eq (3.5) in chapter 3 of [lohr2021sampling] and the definition of variance of the gradients $(\sigma_{k}^{(g),e-1})^{2}$ evaluated at the particular local step e1e-1 for the parameter realization $\bm{w}_{k}^{(g),e-1}$. Expanding $(\sigma_{k}^{(g),e-1})^{2}$ yields
$(\sigma_{k}^{(g),e-1})^{2}$
where $d$ denotes the feature vector of data point $d$, and $\bm{\lambda}_{k}^{(g)}$ and $(\widetilde{\sigma}_{k}^{(g)})^{2}$ denote the mean and sample variance of data points in $\mathcal{D}_{k}^{(g)}$, which are gradient independent. Also, in (i) of (B), we used the Cauchy-Schwarz inequality, and in (ii), we used the fact that $\sum_{d\in\mathcal{D}_{k}^{(g)}}\left(d-\bm{\lambda}_{k}^{(g)}\right)=0.$ Plugging (B) into (B) gives
Using this, (B) becomes
Rearranging the terms, we have
$\mathbb{E}_{g}\big[F^{(g)}(\bm{w}^{(g+1)})\big]$
Term (b) in (B) can be upper bounded as
$\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
The last term in (57) can be further upper bounded
Combining (58) with (57) gives
$\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
Taking the summation of both sides in (59) across local training steps
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
After some simplification, we obtain
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
Assuming
$\eta^{(g)}\leq\Big(2\beta^{2}\sqrt{e_{k}^{(g)}(e_{k}^{(g)}-1)}\Big)^{-1},$
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
Plugging (62) into (B) gives
To eliminate term (c) in (63), we need to make
we require
Also, we aim to make
$1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)>0.$ (67)
Thus, we get the condition:
Assume there exists a set of constants $\Lambda^{(g)}$, $\forall g$, where:
Then we will have
and
$\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}$
$\frac{1}{G}\sum_{g=1}^{G}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}$
Appendix C Experimental Setup, Reproducibility, and Computational Resources
C-A Setup and Reproducibility
@@
C-B How hyperparameters were chosen
The parameter num_sessions must be at least two greater than num_sessions_pilot to observe the effect of our proposed scheme. The parameters num_rounds_pilot, num_rounds_actual, num_SGD_training, num_SGD_grad_cal, and num_round_grad_cal all take positive integer values, while similarity_scale is a positive real number. We experimented with values such as 10, 100, 500, and 800, and selected appropriate settings based on the model and dataset. In principle, most hyperparameters were chosen such that training within each session converges by its final round, using the smallest values possible to ensure fair comparison across algorithms and datasets.
Algorithm-specific parameters include prox_alpha, moon_mu, moon_tau, kl_coefficient, acg_beta, and acg_lambda. These were selected based on the model architecture, data distribution, and dataset characteristics, or by following the recommended values from the respective original papers.
C-C Baseline variants:
AG News uses the average and previous baselines. All other datasets use average, previous, and continuous baselines.
@@ C-D Compute Resources
Table LABEL:table:resource_20_appendix presents a comprehensive overview of average CPU time and maximum memory usage for our experiments with 20 clients. The average is taken over three random seeds.
TABLE VIII: Average CPU time and average maximum memory usage across three random seeds for our experiments with 20 clients.
@@
Table LABEL:table:resource_100_appendix presents a comprehensive overview of average CPU time and average maximum memory usage for our experiments with 100 clients. The average is taken over three random seeds.
TABLE IX: Average CPU time and average maximum memory usage across three random seeds for our experiments with 100 clients.
@@
Appendix D Supplementary Results on Label Distributions from the Main Text
Tables X further demonstrates the effectiveness of Algorithm 1 on the MNIST, Fashion-MNIST, SVHN, CIFAR-10, and CIFAR-100 datasets with 100 clients.
TABLE X: Performance comparison of implementing Algorithm 1 with MOON, FedProx, SCAFFOLD and FedACG under different label distributions for MNIST, Fashion-MNIST (FMNIST), SVHN, and CIFAR-10 dataset for 100 clients.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. Second Session Transition Fourth Session Transition
CIFAR-100 0.0 0.3 MOON 38.74±0.52 32.35±0.52 37.40±0.61 36.09±0.58 42.21±0.55 38.76±0.46 43.42±0.85 39.70±0.49
FedProx 38.38±0.46 32.21±0.46 35.87±1.13 35.41±0.56 41.94±0.40 38.83±0.68 42.04±1.14 39.23±0.54
SCAFFOLD 45.03±0.34 46.10±0.80 44.21±0.65 42.32±0.56 48.32±0.76 50.89±0.50 49.94±0.34 45.49±0.50
FedACG 47.07±0.37 32.30±0.19 26.86±0.57 41.86±0.80 50.10±0.46 38.90±0.36 38.18±0.86 45.26±0.55
0.7 MOON 41.25±0.65 35.54±0.52 39.22±1.20 38.10±0.64 44.99±0.89 41.91±0.49 44.65±1.87 41.43±0.54
FedProx 41.11±0.36 35.62±0.53 37.07±1.67 37.70±0.65 44.59±0.54 42.02±0.49 42.71±1.32 41.60±0.72
SCAFFOLD 47.60±0.43 47.85±0.46 46.91±0.58 44.16±0.59 50.71±0.70 51.37±0.44 52.50±0.81 47.79±0.42
FedACG 49.57±0.40 35.58±0.48 28.44±0.62 44.35±0.94 52.30±0.78 42.13±0.41 39.32±1.57 47.63±0.51
0.2 0.3 MOON 44.46±0.65 39.95±0.81 44.49±0.64 35.92±1.05 47.35±0.75 45.93±0.27 49.99±0.81 39.61±0.47
FedProx 44.18±1.00 40.27±0.38 45.57±0.99 38.69±0.36 47.40±0.90 45.92±0.28 50.22±0.34 42.32±0.47
SCAFFOLD 49.40±0.53 49.20±0.90 47.35±1.00 43.93±0.43 52.22±0.38 51.88±0.55 52.19±1.35 47.31±0.41
FedACG 51.73±0.22 39.95±0.83 39.87±1.36 43.93±1.12 53.64±0.19 45.62±0.70 46.71±0.92 48.03±1.22
0.7 MOON 47.27±0.41 43.14±0.31 44.98±0.29 37.88±0.66 50.41±0.33 48.62±0.40 51.39±0.64 41.56±0.58
FedProx 46.87±0.33 43.05±0.51 47.72±0.19 40.45±0.30 49.57±0.52 48.67±0.38 52.18±0.29 44.14±0.52
SCAFFOLD 51.28±0.74 51.47±0.46 51.13±0.39 47.94±0.47 54.08±0.94 54.61±0.38 56.20±0.20 51.45±0.53
FedACG 53.52±0.14 42.91±0.44 41.16±1.25 47.39±0.54 55.82±0.26 48.58±0.29 48.99±0.75 51.11±0.48
Appendix E Additional Results for Alternative Label Distributions
  • Two-Shard [mcmahan2017communication, hsu2019measuring, li2020federated, fallah2020personalized, karimireddy2020scaffold]: For datasets with 10 labels, each label is divided into two equally-sized shards, and each client receives two different label shards. For datasets with 100 labels, the labels are divided into 10 non-overlapping batches, and each batch is split into two shards, which are then assigned to two randomly selected clients. Each client ends up with data from two labels for 10-label datasets, or 20 labels for 100-label datasets .
Table XI compares the average accuracy of the proposed algorithm across all FL algorithms during the first 10 global rounds after three data distribution shifts under the “Two-Shard” and “Partial-Overlap” label distributions. The results across datasets, label distributions, and models show that the proposed algorithm mitigates performance degradation from dynamic client arrivals and departures. Moreover, it accelerates performance recovery when the current data distribution aligns with a previously observed one.
TABLE XI: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and MOON under different label distributions and datasets. Performance is measured across 3 transitions for each dataset.
SVHN (CNN) 86.52 40.13 72.7 97.44 84.53 95.08 95.8 55.16 83.08
SVHN (CNN) 92.4 76.09 91.72 93.64 68.67 92.35 92.51 71.79 92.02
SVHN (CNN) 76.53 51.03 56.65 83.45 39.58 72.65 95.87 42.97 85.15
SVHN (CNN) 92.27 79.11 89.94 93.67 82.01 91.46 92.48 77.2 91.47
SVHN (CNN) 79.23 69.2 47.94 95.32 90.76 73.11 93.96 88.37 73.25
CIFAR100 (ResNet18) 21.63 5.61 9.59 35.75 21.3 19.42 26.18 7.69 11.98
SVHN (CNN) 90.33 84.02 73.88 91.9 89.05 80.91 91.33 85.83 74.68
SVHN (CNN) 87.0 78.61 75.72 69.51 64.57 54.88 93.49 90.2 89.87
SVHN (CNN) 92.96 90.41 88.39 91.74 87.97 90.77 93.18 91.25 90.25
TABLE XII: Performance comparison of FedProx, FedAvg, SCAFFOLD, and MOON under “Half” and “Distinct” label distributions across various datasets. Performance is evaluated over three transitions for each dataset.
SVHN (CNN) 93.23 27.0 92.18 91.64 17.15 90.98 93.53 69.5 92.75
Fashion-MNIST (MLP) 95.5 92.4 95.0 95.8 90.59 94.56 94.02 87.36 90.73
SVHN (CNN) 92.25 45.18 67.51 87.91 38.45 63.85 92.24 42.7 60.68
SVHN (CNN) 91.63 71.49 88.78 93.54 51.98 91.07 91.98 42.34 90.55
SVHN (CNN) 94.99 49.88 80.62 91.9 38.94 58.53 87.21 32.71 66.47
SVHN (CNN) 91.66 86.87 76.06 89.86 84.79 80.25 93.04 87.31 87.18
SVHN (CNN) 93.64 89.35 89.95 92.11 87.3 91.18 93.63 90.54 91.37
Table XII presents a comparative analysis of the average accuracy achieved by the proposed algorithm across all federated learning (FL) algorithms during the first 10 global rounds following three shifts in data distribution under the “Half” and “Distinct” label distributions.
Table XIII presents the results of applying Table 1 with FedAvg for 20 clients on the TinyImageNet dataset [le2015tiny]. The baseline method used is “Average.” These results highlight the effectiveness of Table 1 in handling scenarios with a larger client population and a more complex dataset using a more advanced model ResNet34.
TABLE XIII: Performance of Algorithm 1 with FedAvg under “Half” and “Partial-Overlap” label distributions for the TinyImageNet dataset with 20 clients. Performance is measured across three transitions for each dataset.
Proposed Previous Proposed Previous Proposed Previous
Half TinyImageNet (ResNet34) 80.72 72.81 77.77 72.24 80.43 73.42
Table XIV shows the results for 100 clients for Algorithm 1 integrated with FedProx, FedAvg, SCAFFOLD, and MOON on the MNIST and Fashion-MNIST datasets, using 100 clients under the Partial-Overlap label distribution. Performance is evaluated over three transitions for each dataset.
TABLE XIV: Performance comparison of Algorithm 1 integrated with FedProx, FedAvg, SCAFFOLD, and MOON on the MNIST and Fashion-MNIST datasets, utilizing 100 clients for FedProx, FedAvg, and MOON, and 50 clients for SCAFFOLD under the Partial-Overlap label distribution. Performance is evaluated across three transitions for each dataset.
MOON MNIST (MLP) 91.36 87.4 90.45 94.33 89.92 90.78 91.55 87.87 90.55